OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 243 and a4 = 9

OpenStudy (binary3i):

Geometric sequence a, ax, ax^2, ax^3, ax^4.............. a=243 ax^3=9 ax^3/a=9/243=1/27=x^3 so x= 1/3 now the 8th term+ ax^7=243(1/3)^7 now you sole it you'll get the answer.

OpenStudy (anonymous):

yeah i got it thankkk you(: